# Towards Prospective Medical Image Reconstruction via Knowledge-Informed Dynamic Optimal Transport

### v1

@@ Abstract
Medical image reconstruction from measurement data is a vital but challenging inverse problem. Deep learning approaches have achieved promising results, but often requires paired measurement and high-quality images, which is typically simulated through a forward model, i.e., retrospective reconstruction. However, training on simulated pairs commonly leads to performance degradation on real prospective data due to the retrospective-to-prospective gap caused by incomplete imaging knowledge in simulation. To address this challenge, this paper introduces imaging Knowledge-Informed Dynamic Optimal Transport (KIDOT), a novel dynamic optimal transport framework with optimality in the sense of preserving consistency with imaging physics in transport, that conceptualizes reconstruction as finding a dynamic transport path. KIDOT learns from unpaired data by modeling reconstruction as a continuous evolution path from measurements to images, guided by an imaging knowledge-informed cost function and transport equation. This dynamic and knowledge-aware approach enhances robustness and better leverages unpaired data while respecting acquisition physics. Theoretically, we demonstrate that KIDOT naturally generalizes dynamic optimal transport, ensuring its mathematical rationale and solution existence. Extensive experiments on MRI and CT reconstruction demonstrate KIDOT’s superior performance.
@@ 1 Introduction
Medical imaging techniques like Magnetic Resonance Imaging (MRI) and Computed Tomography (CT) are vital tools for visualizing internal anatomy, yet reconstructing high-fidelity images from the acquired measurements remains a significant challenge [1, 2]. This process is fundamentally an ill-posed inverse problem inferring a complete, clean image from often incomplete and noisy measurement data. Classical reconstruction algorithms, such as filtered back-projection or iterative methods [3, 4, 5, 6, 7], typically rely on precise mathematical modeling of the image acquisition physics, incorporating essential imaging knowledge. However, accurately capturing complex factors present in real-world scenarios (e.g., noise, non-ideal sampling) within these models remains challenging. Consequently, the performance of traditional methods can degrade significantly when this knowledge is incomplete or imperfect, yielding images with noise or residual artifacts.
Deep learning (DL) has emerged as a powerful alternative [8, 9, 10], achieving state-of-the-art results by learning complex mappings directly from data. However, the predominant supervised DL paradigm[11, 12, 13] faces a critical bottleneck: the requirement for large datasets of perfectly registered pairs of low-quality input measurements and their corresponding high-quality ground truth images. Obtaining such perfectly aligned data in clinical practice is often difficult, costly, or ethically constrained. Furthermore, many real-world datasets inherently lack perfect pairing due to patient motion, or differences in acquisition protocols between low-quality and high-quality scans [14, 15].
@@
To circumvent the lack of real paired data, a common practice [16, 17] is to generate simulated low-quality data using simplified imaging knowledge (e.g., adding modeled noise, applying idealized undersampling masks) from available high-quality images, creating artificial pairs for supervised training. However, a significant gap often exists between these simulations and real-world prospective measurements. Figure 1(b) shows examples of the retrospective-prospective gap in MRI, where k-space data from retrospective simulations (left) differ from that of prospective acquisitions (right) in sampling patterns and reconstructed high-frequency regions. Retrospective simulations often idealize the true data acquisition process, struggling to fully capture the complex noise and genuine artifact characteristics inherent in real-world prospective measurements. Models trained solely on such simulated pairs, therefore, often exhibit a performance drop when deployed in prospective clinical settings, highlighting their sensitivity to distribution shifts arising from the disparity between simulated and real-world imaging knowledge. Ideally, a reconstruction method should effectively leverage all available data resources, including potentially large amounts of unpaired real measurements and real high-quality images, while robustly incorporating reliable imaging knowledge. This motivates exploring frameworks capable of learning from unpaired distributions in a physics-aware manner. Optimal Transport (OT) [18, 19] offers a principled mathematical approach for transporting probability distributions or finding mappings between them, with the potential to learn from unpaired data. However, conventional applications of OT often adopt a static perspective, learning a direct map and potentially overlooking the dynamic physical processes inherent in image reconstruction.
@@ 2 Background
Medical imaging seeks to reconstruct internal structures from indirect, often noisy measurements. This task is fundamentally an ill-posed inverse problem: determining an underlying image $x$ given observed data $y$. The relationship is typically modeled as $y=\mathcal{A}(x)+n$, where $\mathcal{A}$ represents the forward model dictated by the imaging physics (e.g., the Radon transform in CT [20] or the partial Fourier transform in MRI [21]), and $n$ accounts for measurement noise and errors. Current approaches to solving such inverse problems can be broadly categorized into three types. Classical iterative methods [3, 4] minimize the difference between predicted and measured data, guided by the physical model $\mathcal{A}$ and incorporating prior assumptions about the image $x$ (e.g., sparsity [3] or low-rank structure [4]) via regularization:
@@
where $\mathcal{R}(x)$ is the regularization term, and $\lambda$ is the trade-off parameter. While grounded in physics, these methods may struggle to recover fine details from limited or noisy data. More recently, deep learning techniques [22, 23, 24, 25, 26] have shown promise, learning direct mappings from measurements $y$ to images $x$. These data-driven models can achieve high accuracy and speed but typically require large datasets of paired examples (i.e., measurement $y$ and corresponding ground truth image $x$). Hybrid approaches [11, 27] attempt to combine the strengths of both data- and model-driven methods, integrating physical models within neural network architectures. However, like purely data-driven methods, they usually depend on the availability of paired training data. A significant practical challenge arises because acquiring such perfectly aligned, high-quality reference images alongside clinical measurements is often difficult or impossible (e.g., obtaining perfectly matched low-dose and standard-dose CT scans). This common scenario, where only measurements $y$ are readily available without corresponding high-quality $x$, defines the challenging prospective inverse problem setting that remains underexplored [16, 17].
@@
OT offers a robust mathematical framework for comparing probability distributions via the minimal cost required to transform one into another [28, 18]. Static OT finds an optimal coupling $\pi$ between source $\mathbb{P}$ and target $\mathbb{Q}$ distributions that minimizes the expected transport cost [18]:
@@
where $c$ is the cost function, and $\Pi$ is the set of joint distributions with marginals $\mathbb{P},\mathbb{Q}$. Extending this, dynamic OT models the continuous evolution from $\mathbb{P}$ to $\mathbb{Q}$ over time [18, 19]. Dynamic OT seeks a velocity field $v(t,x(t))$ that optimally transports particles from an initial distribution $x(0)\sim\mathbb{P}$ to a final distribution $x(1)\sim\mathbb{Q}$. This is achieved by finding $v$ that minimizes the action functional:
@@
The OT principles have proven valuable for inverse problems lacking paired data, e.g., natural image processing [29, 30, 31, 32, 33]. In medical imaging, static OT has been explored to address the challenges of unpaired or misaligned clinical data. For example, [34, 35] use static OT to derive a CycleGAN framework for learning from unpaired data. [36] further integrates variational methods with reconstruction and Wasserstein-1 distance for unpaired learning in CT. [37, 38] preserve structural consistency [37] or incorporate contextual features [38] in OT-based generative models. Static OT has also been utilized for aligning and synthesizing different MRI sequences to enhance reconstruction quality [39].
@@ 3.2 Transport Equation Guided by Imaging Knowledge
[…]
@@ 3.3 Practical Implementation Based on Neural Networks
For convenience, we directly parameterize $\nabla\mathcal{R}$ using a neural network $\mathcal{H}_{\phi}$ with parameters $\phi$. To solve problem (7), another challenge is the ODE-based constraint with initial and ending distribution constraints, making optimization difficult. To tackle this challenge, we apply a relaxation to the ending distribution to seek an approximation of the optimal solution. Specifically, given an initial state $I_{0}\sim\mathbb{P}$, we produce $I_{t}$ using ODE by $I_{t}\triangleq\mathcal{T}^{t}_{\phi}(I_{0})=I_{0}-\int_{0}^{t}[\mathcal{A}^{*}(\mathcal{A}(I_{\tau})-I_{0})+\mathcal{H}_{\phi}(I_{\tau})]\mathrm{d}\tau$. By denoting the distribution of $\mathcal{T}^{1}_{\phi}(I_{0})$ as $\mathcal{T}^{1}_{\phi\#}(\mathbb{P})$, the constraint $I_{1}\sim\mathbb{Q}$ becomes $\mathcal{T}^{1}_{\phi\#}(\mathbb{P})=\mathbb{Q}$, equivalent to $W_{1}(\mathcal{T}^{1}_{\phi\#}(\mathbb{P}),\mathbb{Q})=0$ where $W_{1}$ is the 1-Wasserstein distance. By introducing the Lagrange relaxation, problem (7) becomes
@@
where $\mathcal{C}(\phi)=\mathbb{E}_{I_{0}\sim\mathbb{P}}\left[\int_{0}^{1}c(\mathcal{T}_{\phi}^{t}(I_{0}),I_{0})\,\mathrm{d}t\right]$. Leveraging the Kantorovich-Rubinstein duality [18] for the $W_{1}$ distance, we further transform the optimization problem to
$\inf_{\phi}\left\{\mathcal{C}(\phi)+\lambda\sup_{\|\varphi\|_{\mathrm{Lip}}\leq 1}\left(\mathbb{E}_{I_{0}\sim\mathbb{P}}[\varphi(\mathcal{T}_{\phi}^{1}(I_{0}))]-\mathbb{E}_{x\sim\mathbb{Q}}[\varphi(x)]\right)\right\},$ (9)
@@
To solve the above problem, we also parameterize $\varphi$ using a neural network $\varphi_{\theta}$ with parameters $\theta$. The integral to produce $\mathcal{T}_{\phi}^{t}(I_{0})$ and the cost are approximated numerically via a discrete sum over $N$ time steps $10=t_{0}<t_{1}<\cdots<t_{N}=$ 1, with a step size $t_{i+1}-t_{i}=1/N$. Using a forward Euler discretization for the ODE in Eq. (7), the evolution at each discrete step becomes $I_{t_{i+1}}=I_{t_{i}}-\frac{1}{N}\left(\mathcal{A}^{*}(\mathcal{A}(I_{t_{i}})-I_{0})+\mathcal{H}_{\phi}(I_{t_{i}})\right)$, where $I_{t_{0}}=I_{0}$. Let $\mathcal{T}_{\phi}^{t_{i}}(I_{0})$ denote the state $I_{t_{i+1}}$ obtained after $i+1$ such steps. Substituting these into the dual objective Eq. (9) and discretizing the time integral, we arrive at the following loss function for training:
@@ 4 Experiments
Our experiments on simulated MRI data utilized the publicly accessible fastMRI multi-coil knee dataset [40]. We selected 2500 fully sampled MRI slices, partitioning them into 1000 for training, 500 for validation, and 1000 for testing. Undersampled measurements were simulated using k-space masks corresponding to a 4x acceleration factor. A key aspect of this evaluation was to assess KIDOT’s robustness to scenarios where test data characteristics differ from training data, a common challenge in prospective clinical deployment. To achieve this, we generated two distinct sets of undersampled inputs: one set, with specific sampling patterns, was used for training supervised baselines, while a different set of sampling patterns was employed for testing to mimic prospective acquisition conditions. For this dataset, the core transport network $\mathcal{T}_{\phi}$ within KIDOT was implemented using an unfolding architecture based on E2E-VarNet [12].
Quantitative results on the simulated MRI dataset are presented in Table 1. KIDOT demonstrates superior performance across all evaluated metrics, encompassing both image fidelity measures (PSNR and SSIM) and perceptual quality indicators (FID [41] and KID [42]). These scores outperform conventional methods like CS-wavelet [43, 44] and the supervised E2E-VarNet [12], in addition to contemporary unpaired learning techniques such as OT-CycleGAN [34] and UAR [36]. This highlights KIDOT’s effectiveness in producing reconstructions that achieve both high fidelity to the ground truth and strong perceptual realism, even under these challenging simulated prospective conditions. Visual comparisons, further illustrating these improvements, are provided in Appendix B.6.
Table 1: Quantitative comparison on the simulated MRI dataset (4x acceleration). Metrics include PSNR, SSIM (higher is better $\uparrow$), FID, and KID×$\times$100 (lower is better $\downarrow$). Best results are bolded.
      Method       PSNR $\uparrow$       SSIM $\uparrow$       FID $\downarrow$       KID×$\times$100 $\downarrow$
      CS-wavelet [43]       29.17       0.7981       95.76       5.77
      E2E-VarNet [12]       32.56       0.8372       24.35       1.29
      OT-CycleGAN [34]       26.45       0.7751       156.43       7.25
      UAR [36]       33.01       0.8425       19.56       1.05
@@
Table 2: Quantitative comparison on the Real Prospective MRI Data (10$\times$ acceleration). Metrics include FID and KID×$\times$100 (lower is better $\downarrow$). Best results are bolded.
Method FID $\downarrow$ KID×$\times$100 $\downarrow$
CS-wavelet [43] 67.75 3.34
E2E-VarNet [12] 65.27 2.83
PromptMR [13] 29.52 1.46
OT-CycleGAN [34] 221.83 10.94
UAR [36] 29.09 1.23
@@
For this dataset, the transport network $\mathcal{T}_{\phi}$ in KIDOT was parameterized using an unfolding architecture based on PromptMR [13]. Given the inherent spatial misalignments, which render pixel-wise metrics unreliable, our evaluation on this dataset focused on distribution-based metrics: FID and KID. As detailed in Table 2, KIDOT achieved the leading FID and KID scores, demonstrating a performance advantage over supervised learning baselines such as E2E-VarNet and PromptMR. These results indicate that KIDOT generates reconstructions that are distributionally closer to the fully sampled reference images compared to the other evaluated techniques. Visual results presented in Figure 3 corroborates these quantitative findings. On representative test slices, KIDOT reconstructions exhibit clear anatomical detail, effective noise suppression, and reduced Gibbs artifacts. Particularly in challenging areas (e.g., highlighted region in red), KIDOT demonstrates superior preservation of fine structures compared to other approaches.
@@
BM3D [45] 100.82 6.12
REDCNN [46] 46.43 1.93
ISTA-Net [47] 24.86 1.59
OT-CycleGAN [34] 63.54 3.28
UAR [36] 23.82 1.45
@@
Clinical prospective CT data.
We finally evaluated KIDOT on a challenging clinical low-dose CT (LDCT) task, utilizing a dataset of prospectively acquired abdominal scans from 30 consented patients under Institutional Review Board approval. The LDCT protocol involved an approximate 10-fold reduction in radiation dose compared to the corresponding normal-dose CT (NDCT) scans. Crucially, differences in scan parameters and acquisition directions between the LDCT and NDCT protocols resulted in significant spatial misalignments between the image pairs, a common complication in prospective clinical studies. From 28,251 available slices, 2,420 were selected for this evaluation. To benchmark KIDOT against methods reliant on aligned data, we generated simulated low-dose CT (SLDCT) images from the NDCT scans, injecting compound Gaussian and Poisson noise into the uncorrupted projections by using the method in [48, 49]. The transport network $\mathcal{T}_{\phi}$ within KIDOT for this task was based on the ISTA-Net architecture [47].
@@
We conduct an ablation study on the fastMRI knee dataset (4x acceleration) to evaluate the effectiveness of each component in KIDOT. We compared four different configurations: (1) training with only supervised loss $\mathcal{L}_{\rm SUP}$ in Eq. (11), (2) training with the loss associated with $\varphi_{\theta}$ in Eq. (10) and $\mathcal{L}_{\rm SUP}$, denoted as $\mathcal{L}_{\rm SUP}+\varphi_{\theta}$, (3) adding the transport cost $\|y-\mathcal{A}(\mathcal{T}_{\phi}^{t_{N-1}}(y))\|_{1}$ for the final output in Eq. (10) to $\mathcal{L}_{\rm SUP}+\varphi_{\theta}$, denoted as $\mathcal{L}_{\rm SUP}+\varphi_{\theta}+C_{\rm final}$, (4) using whole loss (KIDOT). Table 4 (upper two rows) shows that adding the potential network $\varphi_{\theta}$ to the supervised baseline resulted in performance improvements. These gains indicate that $\varphi_{\theta}$ successfully guides the learning process towards better alignment with the target data distribution. It can be observed from Table 4 that incorporating the imaging-knowledge-informed transport cost to the final output can further improve performance. Furthermore, our full KIDOT model that enforces consistency to imaging knowledge throughout the transport process, further improves the results, demonstrating the effectiveness of our imaging knowledge-informed dynamic OT for medical image reconstruction.

### v3

@@ Abstract
Medical image reconstruction from measurement data is a vital but challenging inverse problem. Deep learning approaches have achieved promising results, but often requires paired measurement and high-quality images, which is typically simulated through a forward model, i.e., retrospective reconstruction. However, training on simulated pairs commonly leads to performance degradation on real prospective data due to the retrospective-to-prospective gap caused by incomplete imaging knowledge in simulation. To address this challenge, this paper introduces imaging Knowledge-Informed Dynamic Optimal Transport (KIDOT), a novel dynamic optimal transport framework with optimality in the sense of preserving consistency with imaging physics in transport, that conceptualizes reconstruction as finding a dynamic transport path. KIDOT learns from unpaired data by modeling reconstruction as a continuous evolution path from measurements to images, guided by an imaging knowledge-informed cost function and transport equation. This dynamic and knowledge-aware approach enhances robustness and better leverages unpaired data while respecting acquisition physics. Theoretically, we demonstrate that KIDOT naturally generalizes dynamic optimal transport, ensuring its mathematical rationale and solution existence. Extensive experiments on MRI and CT reconstruction demonstrate KIDOT’s superior performance. Code is available at https://github.com/TaoranZheng717/KIDOT.
@@ 1 Introduction
Medical imaging techniques like Magnetic Resonance Imaging (MRI) and Computed Tomography (CT) are vital tools for visualizing internal anatomy, yet reconstructing high-fidelity images from the acquired measurements remains a significant challenge [najjar2023redefining, hussain2022modern]. This process is fundamentally an ill-posed inverse problem inferring a complete, clean image from often incomplete and noisy measurement data. Classical reconstruction algorithms, such as filtered back-projection or iterative methods [daubechies2004iterative, jin2016general, vandenberghe2001iterative, beister2012iterative, jiang2003convergence], typically rely on precise mathematical modeling of the image acquisition physics, incorporating essential imaging knowledge. However, accurately capturing complex factors present in real-world scenarios (e.g., noise, non-ideal sampling) within these models remains challenging. Consequently, the performance of traditional methods can degrade significantly when this knowledge is incomplete or imperfect, yielding images with noise or residual artifacts.
Deep learning (DL) has emerged as a powerful alternative [sarker2021deep, dargan2020survey, janiesch2021machine], achieving state-of-the-art results by learning complex mappings directly from data. However, the predominant supervised DL paradigm[yang2018admm, sriram2020end, PromptMR] faces a critical bottleneck: the requirement for large datasets of perfectly registered pairs of low-quality input measurements and their corresponding high-quality ground truth images. Obtaining such perfectly aligned data in clinical practice is often difficult, costly, or ethically constrained. Furthermore, many real-world datasets inherently lack perfect pairing due to patient motion, or differences in acquisition protocols between low-quality and high-quality scans [tajbakhsh2020embracing, hu2023image].
@@
To circumvent the lack of real paired data, a common practice [almansour2022deep, xie2024prospective] is to generate simulated low-quality data using simplified imaging knowledge (e.g., adding modeled noise, applying idealized undersampling masks) from available high-quality images, creating artificial pairs for supervised training. However, a significant gap often exists between these simulations and real-world prospective measurements. Figure 1(b) shows examples of the retrospective-prospective gap in MRI, where k-space data from retrospective simulations (left) differ from that of prospective acquisitions (right) in sampling patterns and reconstructed high-frequency regions. Retrospective simulations often idealize the true data acquisition process, struggling to fully capture the complex noise and genuine artifact characteristics inherent in real-world prospective measurements. Models trained solely on such simulated pairs, therefore, often exhibit a performance drop when deployed in prospective clinical settings, highlighting their sensitivity to distribution shifts arising from the disparity between simulated and real-world imaging knowledge. Ideally, a reconstruction method should effectively leverage all available data resources, including potentially large amounts of unpaired real measurements and real high-quality images, while robustly incorporating reliable imaging knowledge. This motivates exploring frameworks capable of learning from unpaired distributions in a physics-aware manner. Optimal Transport (OT) [villani2008optimal, chen2021stochastic] offers a principled mathematical approach for transporting probability distributions or finding mappings between them, with the potential to learn from unpaired data. However, conventional applications of OT often adopt a static perspective, learning a direct map and potentially overlooking the dynamic physical processes inherent in image reconstruction.
@@ 2 Background
Medical imaging seeks to reconstruct internal structures from indirect, often noisy measurements. This task is fundamentally an ill-posed inverse problem: determining an underlying image $x$ given observed data $y$. The relationship is typically modeled as $y=\mathcal{A}(x)+n$, where $\mathcal{A}$ represents the forward model dictated by the imaging physics (e.g., the Radon transform in CT [buzug2011computed] or the partial Fourier transform in MRI [vlaardingerbroek2013magnetic]), and $n$ accounts for measurement noise and errors. Current approaches to solving such inverse problems can be broadly categorized into three types. Classical iterative methods [daubechies2004iterative, jin2016general] minimize the difference between predicted and measured data, guided by the physical model $\mathcal{A}$ and incorporating prior assumptions about the image $x$ (e.g., sparsity [daubechies2004iterative] or low-rank structure [jin2016general]) via regularization:
@@
where $\mathcal{R}(x)$ is the regularization term, and $\lambda$ is the trade-off parameter. While grounded in physics, these methods may struggle to recover fine details from limited or noisy data. More recently, deep learning techniques [mardani2018deep, lee2017deep, ahishakiye2021survey, zhang2020review, wang2020deep] have shown promise, learning direct mappings from measurements $y$ to images $x$. These data-driven models can achieve high accuracy and speed but typically require large datasets of paired examples (i.e., measurement $y$ and corresponding ground truth image $x$). Hybrid approaches [yang2018admm, zhang2018ista] attempt to combine the strengths of both data- and model-driven methods, integrating physical models within neural network architectures. However, like purely data-driven methods, they usually depend on the availability of paired training data. A significant practical challenge arises because acquiring such perfectly aligned, high-quality reference images alongside clinical measurements is often difficult or impossible (e.g., obtaining perfectly matched low-dose and standard-dose CT scans). This common scenario, where only measurements $y$ are readily available without corresponding high-quality $x$, defines the challenging prospective inverse problem setting that remains underexplored [almansour2022deep, xie2024prospective].
@@
OT offers a robust mathematical framework for comparing probability distributions via the minimal cost required to transform one into another [villani2008optimal, wang2025joint, gu2022keypoint]. Static OT finds an optimal coupling $\pi$ between source $\mathbb{P}$ and target $\mathbb{Q}$ distributions that minimizes the expected transport cost [villani2008optimal]:
@@
where $c$ is the cost function, and $\Pi$ is the set of joint distributions with marginals $\mathbb{P},\mathbb{Q}$. Extending this, dynamic OT models the continuous evolution from $\mathbb{P}$ to $\mathbb{Q}$ over time [villani2008optimal, chen2021stochastic]. Dynamic OT seeks a velocity field $v(t,x(t))$ that optimally transports particles from an initial distribution $x(0)\sim\mathbb{P}$ to a final distribution $x(1)\sim\mathbb{Q}$. This is achieved by finding $v$ that minimizes the action functional:
@@
The OT principles have proven valuable for inverse problems lacking paired data, e.g., natural image processing [gu2023optimal, korotin2023kernel, korotin2023neural, tang2024residualconditioned, tang2025degradation]. In medical imaging, static OT has been explored to address the challenges of unpaired or misaligned clinical data. For example, [sim2020optimal, oh2020unpaired] use static OT to derive a CycleGAN framework for learning from unpaired data. [mukherjee2021end] further integrates variational methods with reconstruction and Wasserstein-1 distance for unpaired learning in CT. [zhu2023optimal, vasa2025context] preserve structural consistency [zhu2023optimal] or incorporate contextual features [vasa2025context] in OT-based generative models. Static OT has also been utilized for aligning and synthesizing different MRI sequences to enhance reconstruction quality [wang2024spatial]. More recently, dynamic OT has been explored for semi-supervised problems, FSBM [theodoropoulos2025feedback] utilizes a distance-preserving objective to build a dynamic transport process in Schrödinger bridge framework.
@@ 3.2 Transport Equation Guided by Imaging Knowledge
$\text{s.t.}\,\frac{\mathrm{d}I_{t}}{\mathrm{d}t}=-(\mathcal{A}^{*}(\mathcal{A}(I_{t})-I_{0})+\nabla\mathcal{R}(I_{t})),\,I_{0}\sim\mathbb{P},I_{1}\sim\mathbb{Q}.$
@@ 3.3 Practical Implementation Based on Neural Networks
For convenience, we directly parameterize $\nabla\mathcal{R}$ using a neural network $\mathcal{H}_{\phi}$ with parameters $\phi$. To solve problem (7), another challenge is the ODE-based constraint with initial and ending distribution constraints, making optimization difficult. To tackle this challenge, we apply a relaxation to the ending distribution to seek an approximation of the optimal solution. Specifically, given an initial state $I_{0}\sim\mathbb{P}$, we produce $I_{t}$ using ODE by $I_{t}\triangleq\mathcal{T}^{t}_{\phi}(I_{0})=I_{0}-\int_{0}^{t}[\mathcal{A}^{*}(\mathcal{A}(I_{\tau})-I_{0})+\mathcal{H}_{\phi}(I_{\tau})]\mathrm{d}\tau$.
By denoting the distribution of $\mathcal{T}^{1}_{\phi}(I_{0})$ as $\mathcal{T}^{1}_{\phi\#}(\mathbb{P})$, the constraint $I_{1}\sim\mathbb{Q}$ becomes $\mathcal{T}^{1}_{\phi\#}(\mathbb{P})=\mathbb{Q}$, equivalent to $W_{1}(\mathcal{T}^{1}_{\phi\#}(\mathbb{P}),\mathbb{Q})=0$ where $W_{1}$ is the 1-Wasserstein distance. By introducing the Lagrange relaxation, problem (7) becomes
@@
where $\mathcal{C}(\phi)=\mathbb{E}_{I_{0}\sim\mathbb{P}}\left[\int_{0}^{1}c(\mathcal{T}_{\phi}^{t}(I_{0}),I_{0})\,\mathrm{d}t\right]$. Leveraging the Kantorovich-Rubinstein duality [villani2008optimal] for the $W_{1}$ distance, we further transform the optimization problem to
$\inf_{\phi}\left\{\mathcal{C}(\phi)+\lambda\sup_{\|\varphi\|_{\mathrm{Lip}}\leq 1}\left(\mathbb{E}_{x\sim\mathbb{Q}}[\varphi(x)]-\mathbb{E}_{I_{0}\sim\mathbb{P}}[\varphi(\mathcal{T}_{\phi}^{1}(I_{0}))]\right)\right\},$ (9)
@@
To solve the above problem, we also parameterize $\varphi$ using a neural network $\varphi_{\theta}$ with parameters $\theta$. The integral to produce $\mathcal{T}_{\phi}^{t}(I_{0})$ and the cost are approximated numerically via a discrete sum over $N$ time steps $0=t_{0}<t_{1}<\cdots<t_{N}=1$, with a step size $t_{i+1}-t_{i}=1/N$. Using a forward Euler discretization for the ODE in Eq. (7), the evolution at each discrete step becomes $I_{t_{i+1}}=I_{t_{i}}-\frac{1}{N}\left(\mathcal{A}^{*}(\mathcal{A}(I_{t_{i}})-I_{0})+\mathcal{H}_{\phi}(I_{t_{i}})\right)$, where $I_{t_{0}}=I_{0}$. Let $\mathcal{T}_{\phi}^{t_{i}}(I_{0})$ denote the state $I_{t_{i+1}}$ obtained after $i+1$ such steps. Substituting these into the dual objective Eq. (9) and discretizing the time integral, we arrive at the following loss function for training:
@@ 4 Experiments
Our experiments on simulated MRI data utilized the publicly accessible fastMRI multi-coil knee dataset [zbontar2018fastmri]. We selected 2500 fully sampled MRI slices, partitioning them into 1000 for training, 500 for validation, and 1000 for testing. Undersampled measurements were simulated using k-space masks corresponding to a 4x acceleration factor. A key aspect of this evaluation was to assess KIDOT’s robustness to scenarios where test data characteristics differ from training data, a common challenge in prospective clinical deployment. To achieve this, we generated two distinct sets of undersampled inputs: one set, with specific sampling patterns, was used for training supervised baselines, while a different set of sampling patterns was employed for testing to mimic prospective acquisition conditions. For this dataset, the core transport network $\mathcal{T}_{\phi}$ within KIDOT was implemented using an unfolding architecture based on E2E-VarNet [sriram2020end].
Quantitative results on the simulated MRI dataset are presented in Table 1. KIDOT demonstrates superior performance across all evaluated metrics, encompassing both image fidelity measures (PSNR and SSIM) and perceptual quality indicators (FID [heusel2017gans] and KID [sutherland2018demystifying]). These scores outperform conventional methods like CS-wavelet [lustig2007sparse, uecker2015berkeley], supervised learning models like E2E-VarNet [sriram2020end], contemporary unpaired learning techniques such as OT-CycleGAN [sim2020optimal], UAR [mukherjee2021end] and FSBM [theodoropoulos2025feedback], as well as strong diffusion-based models like DDS [chung2024decomposed]. This highlights KIDOT’s effectiveness in producing reconstructions that achieve both high fidelity to the ground truth and strong perceptual realism, even under these challenging simulated prospective conditions. Visual comparisons, further illustrating these improvements, are provided in Appendix B.8.
[…]
   CS-wavelet [lustig2007sparse]    29.17    0.7981    95.76    5.77
   E2E-VarNet [sriram2020end]    32.56    0.8372    24.35    1.29
   OT-CycleGAN [sim2020optimal]    26.45    0.7751    156.43    7.25
   UAR [mukherjee2021end]    33.01    0.8425    19.56    1.05
   FSBM [theodoropoulos2025feedback]    26.52    0.7634    113.62    6.03
   DDS [chung2024decomposed]    31.27    0.8136    29.46    1.65
@@
[…]
CS-wavelet [lustig2007sparse] 67.75 3.34
E2E-VarNet [sriram2020end] 65.27 2.83
PromptMR [PromptMR] 29.52 1.46
OT-CycleGAN [sim2020optimal] 221.83 10.94
UAR [mukherjee2021end] 29.09 1.23
FSBM [theodoropoulos2025feedback] 147.62 7.13
DDS (Pre-trained) [chung2024decomposed] 62.94 2.71
DDS (From scratch) [chung2024decomposed] 58.32 2.53
DDS (Fine-tuned) [chung2024decomposed] 39.78 1.84
@@
For this dataset, the transport network $\mathcal{T}_{\phi}$ in KIDOT was parameterized using an unfolding architecture based on PromptMR [PromptMR]. Given the inherent spatial misalignments, which render pixel-wise metrics unreliable, our evaluation on this dataset focused on distribution-based metrics: FID and KID. Furthermore, we conducted a rigorous benchmark against a state-of-the-art DDS model under three practical strategies: using a pre-trained model, fine-tuning it on our data, and training it from scratch. As detailed in Table 2, KIDOT achieved the leading FID and KID scores, outperforming both supervised and unsupervised baselines. KIDOT surpasses all DDS variants, which highlights a key advantage: its physics-informed dynamics provide a powerful inductive bias, making it more data-efficient and robust, especially when real training data is limited. Visual results presented in Figure 3 corroborates these quantitative findings. On representative test slices, KIDOT reconstructions exhibit clear anatomical detail, effective noise suppression, and reduced Gibbs artifacts. Particularly in challenging areas (e.g., highlighted region in red), KIDOT demonstrates superior preservation of fine structures compared to other approaches. Notably, as FSBM is not tailored for medical imaging and its visual results were poor, we omitted its visualizations for brevity.
@@
Clinical prospective CT data.
We finally evaluated KIDOT on a challenging clinical low-dose CT (LDCT) task, utilizing a dataset of prospectively acquired abdominal scans from 30 consented patients under Institutional Review Board approval. The LDCT protocol involved an approximate 10-fold reduction in radiation dose compared to the corresponding normal-dose CT (NDCT) scans. Crucially, differences in scan parameters and acquisition directions between the LDCT and NDCT protocols resulted in significant spatial misalignments between the image pairs, a common complication in prospective clinical studies. From 28,251 available slices, 2,420 were selected for this evaluation. To benchmark KIDOT against methods reliant on aligned data, we generated simulated low-dose CT (SLDCT) images from the NDCT scans, injecting compound Gaussian and Poisson noise into the uncorrupted projections by using the method in [zeng2015simple, li2023noise]. The transport network $\mathcal{T}_{\phi}$ within KIDOT for this task was based on the ISTA-Net architecture [lei2023deep].
@@
BM3D [dabov2007image] 100.82 6.12
REDCNN [chen2017low] 46.43 1.93
ISTA-Net [lei2023deep] 24.86 1.59
OT-CycleGAN [sim2020optimal] 63.54 3.28
UAR [mukherjee2021end] 23.82 1.45
FSBM [theodoropoulos2025feedback] 68.42 3.61
@@
We conduct an ablation study on the fastMRI knee dataset (4x acceleration) to evaluate the effectiveness of each component in KIDOT. We compared four different configurations: (1) training with only supervised loss $\mathcal{L}_{\rm SUP}$ in Eq. (11), (2) training with the loss associated with $\varphi_{\theta}$ in Eq. (10) and $\mathcal{L}_{\rm SUP}$, denoted as $\mathcal{L}_{\rm SUP}+\varphi_{\theta}$, (3) adding the transport cost $C_{\rm final}=\|y-\mathcal{A}(\mathcal{T}_{\phi}^{t_{N-1}}(y))\|_{1}$ for the final output in Eq. (10) to $\mathcal{L}_{\rm SUP}+\varphi_{\theta}$, denoted as $\mathcal{L}_{\rm SUP}+\varphi_{\theta}+C_{\rm final}$, (4) using whole loss (KIDOT). Table 4 (upper two rows) shows that adding the potential network $\varphi_{\theta}$ to the supervised baseline resulted in performance improvements. These gains indicate that $\varphi_{\theta}$ successfully guides the learning process towards better alignment with the target data distribution. It can be observed from Table 4 that incorporating the imaging-knowledge-informed transport cost to the final output can further improve performance. Furthermore, our full KIDOT model that enforces consistency to imaging knowledge throughout the transport process, further improves the results, demonstrating the effectiveness of our imaging knowledge-informed dynamic OT for medical image reconstruction. The key difference between (3) and (4) lies in when and how the imaging- knowledge-informed cost is applied. The third setting represents an ablation where the physics-consistency cost is applied only to the final output of the N-step transport path. In contrast, our full KIDOT model enforces this physics consistency throughout the entire dynamic transport process, integrating the cost over all intermediate steps of the trajectory.